### v2

Abstract
We derive error bounds for CUR matrix approximation using determinant-based methods that relate local projection errors to global approximation quality. For general matrices, we establish determinant identities for bordered Gramian matrices that decompose CUR approximation errors into interpretable local components. These identities connect projection errors onto submatrix column spaces directly to determinants, providing geometric insight into the degradation of the approximation. We develop a probabilistic framework based on volume sampling that yields interpolation-type error bounds quantifying the benefits of oversampling: when $r>k$ rows are sampled to construct a rank-$k$ approximation, the expected error factor decreases linearly from $(k+1)^{2}$ (no oversampling) to $(k+1)$ (full oversampling). Our analysis establishes that the expected squared error is bounded by this interpolation factor times the squared error of the best rank-$k$ approximation, directly connecting the quality of the CUR approximation to that of the optimal low-rank approximation. The framework applies to both CUR decomposition for general matrices and the Nyström method for symmetric positive semi-definite matrices, providing a unified theoretical foundation for the analysis of determinant-based low-rank approximation.
keywords:
Matrix approximation , CUR decomposition , determinants , volume sampling , oversampling
1 Introduction
1.1 Low-Rank Matrix Approximation and the CUR Framework
Low-rank matrix approximation is a fundamental computational primitive in modern data science, underpinning diverse applications ranging from recommendation systems and image compression [2] to kernel methods via the Nyström approach [26], hierarchical solvers for PDEs [3], reduced-order modelling in computational engineering [4], data compression in network monitoring [25], and large-scale machine learning [22].
Given a large matrix $M\in\mathbb{R}^{m\times n}$, the central goal is to find a rank-$k$ approximation with $k\ll\min(m,n)$ that captures the essential structure of the data while dramatically reducing storage and computational requirements.
The classical solution is the truncated singular value decomposition (SVD), which provides the optimal rank-$k$ approximation in all Schatten $p$-norms—a fundamental result established by Eckart and Young [10]. This optimality has made the truncated SVD the gold standard for low-rank approximation [12]. However, computing truncated SVDs requires access to the entire matrix and incurs computational costs that become prohibitive for massive datasets. Moreover, the resulting singular vectors are abstract linear combinations of the original data elements, making them difficult to interpret in many practical applications [21].
An alternative paradigm seeks to approximate matrices directly from subsets of their rows and columns. In its simplest form, one selects $k$ rows and $k$ columns to form an intersection submatrix $A=M_{I,J}$. If $A$ has full rank, the optimal reconstruction of this type takes the form
$M\approx CUR,\quad C=M_{:,J},\;R=M_{I,:},\;U=A^{-1},$
where $A^{-1}$ denotes the inverse of $A$. This construction, known as a skeleton or pseudo-skeleton approximation [14], has the crucial property that it interpolates the exact entries of $M$ in the chosen rows and columns.
In the more general CUR decomposition with oversampling, more than $k$ rows or columns may be selected (i.e., $r>k$) to improve robustness [23]. The theoretically optimal choice for the middle factor is
$U^{\star}=C^{+}MR^{+},$ (1)
as demonstrated by Stewart [24] and employed in the general CUR framework [21]. This choice minimizes the Frobenius norm error over all possible $U$. However, computing $U^{\star}$ requires access to the entire matrix $M$, which defeats the purpose of the submatrix approach. A widely used and computationally efficient alternative retains the simple choice $U=A^{+}$ (the pseudoinverse of the rectangular submatrix), for which sharp deterministic error bounds can still be established. This fundamental trade-off between the optimal but expensive choice $U^{\star}$ and the computationally tractable $A^{+}$ is central to the theoretical development and practical algorithms studied in this paper, particularly in the oversampling regime where $r>k$.
1.2 Determinant Identities and Volume Sampling
The principal contribution of this paper is a comprehensive analysis of CUR and Nyström approximation errors in the Frobenius norm, achieved by relating global error bounds to local quantities expressed through determinants of Gramian matrices. Our approach builds on the observation that determinants naturally encode geometric information about matrix approximations. For instance, when augmenting a submatrix $A$ with an additional column $b$, we have the fundamental identity [19]
$\operatorname{det}([A,b]^{\mathsf{T}}[A,b])=\operatorname{det}(A^{\mathsf{T}}A)\,\|(I-AA^{+})b\|^{2}.$
This identity reveals that local determinants directly encode projection errors—the right-hand side explicitly separates the volume contribution of the existing subspace from the residual component orthogonal to it. Similar formulas hold when adding rows, or simultaneously adding both a row and a column, and these can be unified using compound matrices and the Cauchy–Binet theorem. These local determinant identities form the fundamental building blocks for the global error analysis developed in subsequent sections. The analysis is termed ’local’ because these determinant identities capture the approximation errors introduced by individual rows and columns from the original matrix, allowing us to understand how each additional data element affects the overall CUR decomposition quality.
A key insight emerging from this determinant-based analysis is a family of interpolation-type error bounds that precisely quantify how oversampling improves approximation quality. When $r=k$ (no oversampling), the expected Frobenius error satisfies a $(k+1)^{2}$-factor bound relative to the best rank-$k$ approximation, recovering the classical result of [13]. As the oversampling parameter $r$ increases, this factor decreases linearly, reaching a $(k+1)$-factor when $r=m$. This demonstrates that oversampling rows and columns not only enhances numerical stability but also yields provably sharper error bounds, consistent with recent deterministic analyses [5] and randomized methods [17]. While the bounds on the Frobenius norm of the CUR error also apply to the Nyström method, they are consistent with bounds previously obtained from the eigenvalue decomposition [11].
The analysis relies on connections between matrix approximation errors and elementary symmetric polynomials in the singular values, established through compound matrix theory (detailed in Section 3).
1.3 Historical Development and Recent Advances
An extensive review of low-rank matrix approximations including applications and algorithms can be found in [20].
The theoretical foundation for CUR and related decompositions has evolved through several distinct phases. The earliest roots lie in classical projection theory, where orthogonal projections $AA^{+}$ are used to approximate matrices of the form $[A,B]$, with error bounds expressed through principal angles [6]. These results were subsequently popularized in the numerical linear algebra community through the influential textbook of Golub and Van Loan [12].
A significant advance came with the introduction of rank-revealing factorizations, particularly RRQR decompositions, which provide deterministic methods for selecting columns that preserve numerical rank and approximate singular values. Key contributions in this area include the work of Gu and Eisenstat [16] and other foundational developments in numerical linear algebra. This development strongly influenced later column subset selection strategies and established connections to condition number bounds.
Concurrently, the emergence of randomized numerical linear algebra brought new perspectives to matrix approximation. Notable contributions include the CUR algorithms of Drineas, Mahoney, and collaborators [9, 21] and the comprehensive survey by Halko, Martinsson, and Tropp [17]. These randomized approaches achieve relative-error guarantees in expectation by sampling according to leverage scores or related probability distributions, offering computational advantages for very large matrices. An error bound for the projection method ($A=C$) is established in [8] where it is shown that there exists a submatrix $A\in\mathbb{R}^{m,k}$ of $M$ satisfying
$\|M-AA^{+}M\|_{F}^{2}\leq(k+1)\,\|M-M_{k}\|_{F}^{2},$
where $M_{k}$ denotes the best rank-$k$ approximation of $M$. This established the $(k+1)$-factor that continues to guide theoretical developments in the field.
Recent developments have seen the rise of deterministic algorithms that achieve comparable error bounds without explicit volume computations, most prominently in the work of Cortinovis and Kressner [5]. Simultaneously, there has been renewed interest in determinantal point processes and volume sampling as principled approaches to row and column selection [7], with connections to fast sampling algorithms and improved approximation guarantees.
Modern algorithmic developments include efficient approximate sampling methods for determinantal point processes [15] and scalable implementations for nonsymmetric cases [18]. Additionally, hybrid approaches that combine deterministic and randomized strategies have emerged to balance computational efficiency with approximation accuracy [1].
1.4 Contributions and Paper Organization
This paper provides a unified theoretical framework that bridges local determinant identities with global approximation bounds for both CUR and Nyström methods. Our main contributions are as follows:
Local Error Analysis: In Section 2, we develop a comprehensive local analysis based on determinant identities for bordered Gramian matrices. By augmenting a chosen submatrix with additional rows and/or columns, we derive explicit formulas that decompose global approximation errors into interpretable local components. Our central result (Proposition 3) establishes that for a bordered matrix
$X=\begin{bmatrix}A&b\\ c^{\mathsf{T}}&d\end{bmatrix}$,
one has the explicit identity
$\operatorname{det}(X^{\mathsf{T}}X)=\operatorname{det}(A^{\mathsf{T}}A+cc^{\mathsf{T}})\,\|(I-AA^{+})b\|^{2}+\operatorname{det}(A^{\mathsf{T}}A)\,(d-c^{\mathsf{T}}A^{+}b)^{2}.$
This formula directly relates $\operatorname{det}(X^{\mathsf{T}}X)$ to squared local error terms, providing geometric insight into how matrix approximations degrade when additional data is incorporated.
Local Deterministic Errors: We establish sharp deterministic bounds that require only average-volume rather than maximal-volume assumptions. Specifically, Proposition 7 shows that
$\|X_{\mathrm{CUR}}-X\|^{2}\leq\frac{(r+1)(k+1)}{r+1-k}\,\lambda_{\min}(X^{\mathsf{T}}X),$
provided the selected submatrix has squared volume at least as large as the average over all possible submatrices of the same size. This represents a significant relaxation of traditional maximal-volume requirements.
Global Probabilistic Analysis: Section 3 develops a probabilistic framework based on volume sampling that connects local determinant identities to global approximation guarantees. We derive interpolation-type error bounds that precisely quantify the benefits of oversampling: the error factor transitions smoothly from $(k+1)^{2}$ when $r=k$ to $(k+1)$ when $r=m$, providing both theoretical insight and practical guidance for algorithm design.
In particular, we establish bounds on the expected squared Frobenius error directly in terms of the singular values of $M$, providing a direct connection between the CUR approximation error and the singular value tail:
$\mathbb{E}\left(\|M-M_{\text{CUR}}\|_{F}^{2}\right)\leq\left(\frac{m-r}{m-k}(k+1)^{2}+\frac{r-k}{m-k}(k+1)\right)\sum_{i=k+1}^{n}\sigma_{i}^{2}.$
Together, these contributions provide a comprehensive understanding of the mechanisms governing CUR and Nyström approximation quality, clarifying the role of determinants, volume sampling, and oversampling in achieving accurate low-rank approximations.
The remainder of the paper is organised as follows. Section 2 presents the local determinant analysis and establishes connections to compound matrices via the Cauchy–Binet theorem. Section 3 develops the global volume sampling framework and derives the main approximation bounds. We conclude with a discussion of the implications for algorithm design and directions for future research.